### v1

# The action of an inverse semigroup on its Stone-Čech compactification

Abstract.
We initiate the study of the Stone-Čech transformation groupoid $\mathcal{G}=\mathcal{S}\ltimes\beta\mathcal{S}$ of an inverse semigroup $\mathcal{S}$. This groupoid occurs as a model for the inverse semigroup Roe algebra $\mathcal{R}_{\mathcal{S}}$ of Lledo and Martinez; in particular, the action of $\mathcal{S}$ on $\beta\mathcal{S}$ is amenable in the sense of Exel and Starling if and only if the reduced C-algebra $\textup{C}^{*}_{r}(\mathcal{S})$ is exact. We prove that the properties of being Hausdorff, principal, and effective are all equivalent for $\mathcal{G}$, and give an algebraic condition on $\mathcal{S}$ equivalent to the Hausdorffness of $\mathcal{G}$. Finally, we show that the Hausdorffness of Exel’s tight groupoid $\mathcal{G}_{\text{tight}}(\mathcal{S})$ is necessary for the Hausdorffness of $\mathcal{G}$.
1. Introduction
footnotetext: 2020 Mathematics Subject Classification Primary: 20M18, 18B40, 46L05; Secondary: 43A65.
Keywords: étale groupoids, inverse semigroups, Stone-Čech compactification.
Part of this work was completed as part of the Master’s thesis of the first author. The second author is funded by NSERC grant RGPIN-2021-03834.
A semigroup $\mathcal{S}$ is called inverse if, for every $s\in\mathcal{S}$, there is a unique $s^{*}\in\mathcal{S}$ with $ss^{*}s=s$ and $s^{*}ss^{*}=s^{*}$. The dynamics and geometry of inverse semigroups have been productively studied in a way that generalizes the discrete group setting. In [lledomartinez], Lledo and Martinez described the uniform Roe algebra $\mathcal{R}_{\mathcal{S}}$ of a countable inverse semigroup $\mathcal{S}$, and showed that this C-algebra was naturally isomorphic to the reduced crossed product $\ell^{\infty}(\mathcal{S})\rtimes_{r}\mathcal{S}$.
In this note, we describe an étale groupoid model for $\ell^{\infty}(\mathcal{S})\rtimes_{r}\mathcal{S}$. Like the group case, this étale groupoid arises from a left-translation action of $\mathcal{S}$ on its Stone-Čech compactification (Theorem 3.1). It follows immediately that Exel’s reduced inverse semigroup C-algebra $\textup{C}^{*}_{r}(\mathcal{S})$ is exact if and only if $\mathcal{S}\ltimes\beta\mathcal{S}$ is amenable as an étale groupoid. Said differently, the following generalization of the discrete group case ([ozawa], Theorem 3) holds:
Theorem 1.
Exel’s inverse semigroup C-algebra $\textup{C}^{*}_{r}(\mathcal{S})$ is exact if and only if the action $\mathcal{S}\curvearrowright\beta\mathcal{S}$ is amenable in the sense of Exel and Starling [starling].
It is well-known that the action of a discrete group $\Gamma$ on $\beta\Gamma$ is free (see for example [NR14, Example 2.14]. In the inverse semigroup setting, one may therefore anticipate that $\mathcal{G}=\mathcal{S}\ltimes\beta\mathcal{S}$ is always principal. Using a finite cover criterion of Exel and Pardo [EP16], we obtain the following generalization of this result.
Theorem 2.
Let $\mathcal{S}$ be an inverse semigroup. Then the following statements are equivalent:
  1. (1)
    $\mathcal{S}\ltimes\beta\mathcal{S}$ is Hausdorff,
  2. (2)
    $\mathcal{S}\ltimes\beta\mathcal{S}$ is principal,
  3. (3)
    every clopen action of $\mathcal{S}$ on a locally compact Hausdorff space has a Hausdorff transformation groupoid.
This generalizes the group case because $\mathcal{S}\ltimes\beta\mathcal{S}$ is always Hausdorff when $\mathcal{S}$ is a group. Using Theorem 2, we give a criterion on $\mathcal{S}$ (Theorem 3.8) which implies $\mathcal{S}\ltimes\beta\mathcal{S}$ is Hausdorff for the class of abstract pseudogroups (Corollary 3.9). Finally, since there are many examples of inverse semigroups with non-Hausdorff tight groupoids, condition (3) of Theorem 2 provides many examples of inverse semigroups for which $\mathcal{S}\ltimes\beta\mathcal{S}$ is not Hausdorff nor principal.
2. Preliminaries
In what follows, $\mathcal{S}$ denotes a countable inverse semigroup. An element $e\in\mathcal{S}$ is called idempotent if $e^{2}=e$. For any $\in\mathcal{S}, ss$$^{*}s$ and $ss^{*}$ are idempotent. Denote by $E_{\mathcal{S}}$ the set of idempotent elements of $\mathcal{S}$. It is a pleasant fact that $E_{\mathcal{S}}$ is a commutative subsemigroup of $\mathcal{S}$, and that $\mathcal{S}$ is a group if and only if $E_{\mathcal{S}}$ is a singleton. If $X$ is a set, then the collection of all partial bijections between subsets of $X$, denoted by $\mathcal{I}(X)$, is an inverse semigroup under the operation of partial composition called the symmetric inverse monoid on $X$.
Inverse semigroups carry a natural partial order: $s\leqslant t$ if and only if $ts^{*}s=s$. Restricted to $E_{\mathcal{S}}$, this order is given by $e\leqslant f$ if and only if $ef=e$.
For any poset $(P,\prec)$ and $A\subseteq P$ we will use the notation
$A^{\prec}:=\{b\in P:a\prec b\text{ for some }a\in A\}.$
In this paper, we only consider the partial orders $\subseteq$ on the power set of $\mathcal{S}$, and $\leqslant$ or $\geqslant$ on $\mathcal{S}$.
We now introduce actions of inverse semigroups on topological spaces and C-algebras. Although this note is only interested in two such actions (one on $\beta\mathcal{S}$, and one on $\ell^{\infty}(\mathcal{S})$), we feel it will be beneficial to introduce the theory in somewhat excessive generality. Let $\mathcal{A}$ be a C-algebra. An action of an inverse semigroup $\mathcal{S}$ on $\mathcal{A}$ is a semigroup homomorphism $\alpha:\mathcal{S}\to\mathcal{I}(\mathcal{A})$ such that, for each $s\in\mathcal{S}$,
  1. (1)
    The domain and codomain of $\alpha_{s}$ is a closed, two-sided ideal in $\mathcal{A}$;
  2. (2)
    The partial bijection $\alpha_{s}$ is a *-homomorphism.
Given an idempotent $e\in E_{\mathcal{S}}$, we set $I_{e}=\text{dom}(\alpha_{e})$. We observe that $\alpha_{s}$ is a *-isomorphism of $I_{s^{*}s}$ onto $I_{ss^{*}}$. The next example, central in [lledomartinez], is also the main C-action we will consider in this note.
Example 2.1.
Let $\mathcal{A}=\ell^{\infty}(\mathcal{S})$. For each $e\in E_{\mathcal{S}}$, define $I_{e}=\{f\in\mathcal{A}:\text{supp}(f)\subseteq e\mathcal{S}\}$. For each $s\in\mathcal{S}$, define $\alpha_{s}:I_{s^{*}s}\to I_{ss^{*}}$ by $\alpha_{s}f](t)=f(s^{*}t)$. The mapping $s\in\mathcal{S}\mapsto\alpha_{s}\in\mathcal{I}(\mathcal{S})$ is an action of $\mathcal{S}$ on $\mathcal{A}$.
The C-algebra $\ell^{\infty}(\mathcal{S})$ is *-isomorphic to $C(\beta\mathcal{S})$, where $\beta\mathcal{S}$ is the Stone-Čech compactification of $\mathcal{S}$. At the level of compact Hausdorff spaces, inverse semigroup actions are described as follows (see [EXEL] for details): an action of $\mathcal{S}$ on a compact Hausdorff space $\mathcal{X}$ is a triple $(\mathcal{S},\mathcal{X},\alpha)$, where $\alpha:\mathcal{S}\to\mathcal{I}(\mathcal{X})$ is a semigroup homomorphism such that, for all $s\in\mathcal{S}$,
  1. (1)
    $\alpha_{s}$ is continuous, for all $s\in\mathcal{S}$;
  2. (2)
    $\text{dom}(\alpha_{s})$ is open, for all $s\in\mathcal{S}$.
[…]
Example 2.2 ([gondek24], Section 3.2).
Let us define an inverse semigroup action of $\mathcal{S}$ on the Stone-Čech compactification $(\beta\mathcal{S},\iota)$, where $\mathcal{S}$ is viewed as a discrete space and $\iota:\mathcal{S}\to\beta\mathcal{S}$ is the canonical embedding. For each $s\in\mathcal{S}$, define a bijection $\lambda_{s}:s^{*}\mathcal{S}\to s\mathcal{S}$ by $\lambda_{s}t=st$ ($t\in s^{*}\mathcal{S}$). By the universal property of $\beta\mathcal{S}$, $\lambda_{s}$ extends to a homeomorphism $\lambda_{s}:\overline{s^{*}\mathcal{S}}\to\overline{s\mathcal{S}}$, where $\overline{s^{*}\mathcal{S}}:=\overline{\iota(s^{*}\mathcal{S})}^{\beta}\subseteq\beta\mathcal{S}$. We observe that the domain and codomain of $\lambda_{s}$ are always a compact, open subsets of $\beta\mathcal{S}$. It is routine to check that the mapping $\mathcal{S}\ni s\mapsto\lambda_{s}\in\mathcal{I}(\beta\mathcal{S})$ is a semigroup homomorphism, and therefore an action, which we shall henceforth refer to by $\mathcal{S}\curvearrowright\beta\mathcal{S}$.
The contravariant action $\hat{\lambda}$ of $\mathcal{S}$ on $C(\beta\mathcal{S})$ is the action in Example 2.1, after applying the identification $j:C(\beta\mathcal{S})\ni f\mapsto f\circ\iota\in\ell^{\infty}(\mathcal{S})$. Indeed, given $f\in I_{s^{*}s}\subseteq\ell^{\infty}(\mathcal{S})$, we have $j(f)](t)=f(\iota(t))=0\Leftrightarrow\iota(t)\not\in\overline{s^{*}\mathcal{S}}\Leftrightarrow t\not\in s^{*}\mathcal{S}$, so that $j(\text{dom}(\hat{\lambda}_{s}))=I_{s^{*}s}$, and
$[\alpha_{s}(j(f))](t)=[j(f)](s^{*}t)=f(\iota(s^{*}t))=f(\lambda_{s^{*}}\iota(t))=[\hat{\lambda}_{s}f](t),$
for all $f\in I_{s^{*}s}$.
Given a C-algebra $\mathcal{A}\subseteq\mathcal{B}(\mathcal{H})$ and an inverse semigroup action $\mathcal{S}\curvearrowright\mathcal{A}$, there is a standard construction for a reduced crossed product C-algebra $\mathcal{A}\rtimes_{r}\mathcal{S}\subseteq\mathcal{B}(\mathcal{H}\otimes\ell^{2}(\mathcal{S}))$, which does not depend on the inclusion $\mathcal{A}\subseteq\mathcal{B}(\mathcal{H})$ (for details, see [EXEL]). We will not work directly with this C-algebra, so we only sketch the construction. For each $s\in\mathcal{S}$, we define the maps $V_{s}:\mathcal{S}\to\mathcal{B}(\ell^{2}(\mathcal{S}))$ and $\pi:\mathcal{A}\to\mathcal{B}(\mathcal{H}\otimes\mathcal{\ell}^{2}(\mathcal{S}))$ by
$V_{s}\delta_{t}=\begin{cases}\delta_{st}&t\in s^{*}\mathcal{S}\\ 0&\text{otherwise}\end{cases}\qquad$ $\text{and}\qquad\pi(a)(\xi\otimes\delta_{t})=\begin{cases}\alpha_{t^{*}}(a)\xi\otimes\delta_{t}&t\in I_{s^{*}s}\\ 0&\text{otherwise}.\end{cases}$
The reduced crossed product is the C-subalgebra of $\mathcal{B}(\mathcal{H}\otimes\ell^{2}(\mathcal{S}))$ defined by
$\mathcal{A}\rtimes_{r}\mathcal{S}:=\textup{C}^{*}(\{1\otimes V_{s}:s\in\mathcal{S}\}\cup\pi(A))\subseteq\mathcal{B}(\mathcal{H}\otimes\ell^{2}(\mathcal{S})).$
We are interested in describing an étale groupoid $\mathcal{G}$ with the property that $\text{C}^{*}_{r}(\mathcal{G})=\ell^{\infty}(\mathcal{S})\rtimes_{r}\mathcal{S}$. In preparation for this, let us establish conventions for étale groupoids in this note. Given a groupoid $\mathcal{G}$, we will denote by $\mathcal{G}^{(2)}\subset\mathcal{G}\times\mathcal{G}$ the composable pairs, and $\mathcal{G}^{(0)}=\{\gamma^{-1}\gamma:\gamma\in\mathcal{G}\}$ the unit space. The source and range maps $\mathbf{d}:\mathcal{G}\to\mathcal{G}^{(0)}$ and $\mathbf{r}:\mathcal{G}\to\mathcal{G}^{(0)}$ (respectively) are defined by $\mathbf{d}(\gamma)=\gamma^{-1}\gamma$ and $\mathbf{r}(\gamma)=\gamma\gamma^{-1}$, for $\gamma\in\mathcal{G}$. Given $u\in\mathcal{G}^{(0)}$, we will write $\mathcal{G}_{u}=\mathbf{d}^{-1}(u)$, $\mathcal{G}^{u}=\mathbf{r}^{-1}(u)$, and $\mathcal{G}^{u}_{u}=\mathcal{G}_{u}\cap\mathcal{G}^{u}$. The isotropy $\textup{Iso}(\mathcal{G})$ is defined by
$\textup{Iso}(\mathcal{G})=\{\gamma\in\mathcal{G}:\textbf{r}(\gamma)=\textbf{d}(\gamma)\}.$
We recall that the groupoid $\mathcal{G}$ is called principal if $\textup{Iso}(\mathcal{G})=\mathcal{G}^{(0)}$.
A topological groupoid is a groupoid $\mathcal{G}$ equipped with a topology under which the multiplication $\mathcal{G}^{(2)}\ni(\alpha,\beta)\mapsto\alpha\beta$ and inverse map $\mathcal{G}\ni\gamma\mapsto\gamma^{-1}$ are continuous. A topological groupoid is called étale if it is locally compact, $\mathcal{G}^{(0)}$ is Hausdorff in the subspace topology inherited from $\mathcal{G}$, and the maps $\mathbf{d}$ and $\mathbf{r}$ are local homeomorphisms. We remind the reader that étale groupoids are not required to be Hausdorff. It is a standard fact that an étale groupoid $\mathcal{G}$ is Hausdorff if and only if $\mathcal{G}^{(0)}$ is closed.
Now let $(\mathcal{S},\mathcal{X},\alpha)$ be an inverse semigroup action, where $\mathcal{X}$ is a locally compact Hausdorff space. We associate an étale groupoid to $(\mathcal{S},\mathcal{X},\alpha)$ as follows (see [EXEL] for details). Set
$\Omega=\{(s,x):x\in D_{s^{*}s}\}\subseteq\mathcal{S}\times\mathcal{X}.$
We define an equivalence relation on $\Omega$ by
$(s,x)\sim(t,y)\Leftrightarrow x=y\text{ and }\exists e\in E_{\mathcal{S}}\text{ with }x\in D_{e}\text{ and }se=te.$
One defines a groupoid structure on $\mathcal{G}:=\Omega/{\sim}$ as follows:
  1. (1)
    $\mathcal{G}^{(2)}=\{([s,x],[t,y])\in\mathcal{G}\times\mathcal{G}:x=\alpha_{t}(y)\}$;
  2. (2)
    $s,\alpha_{t}(x)][t,x]=[st,x]$;
  3. (3)
    $s,x]^{-1}=[s^{*},\alpha_{s}(x)]$.
Equipped with the quotient topology, $\mathcal{G}$ becomes an étale groupoid, with basic open neighbourhoods of the form $\Theta(s,U)=\{[s,x]:x\in U\}$ (for $s\in\mathcal{S}$ and open $U\subseteq\mathcal{X}$). We write $\mathcal{G}=\mathcal{S}\ltimes_{\alpha}\mathcal{X}$, and call $\mathcal{G}$ the transformation groupoid (or groupoid of germs) of the action $\mathcal{S}\curvearrowright_{\alpha}\mathcal{X}$.
Every étale groupoid $\mathcal{G}$ generates a reduced C-algebra, $\textup{C}^{*}_{r}(\mathcal{G})$. As we will not have need to directly work with this C-algebra, we refer the reader to [CEPSS] for details. The relation between the two C-algebra constructions described in this section is the content of the next theorem, which is an immediate consequence of ([bussexel], Theorem 4.11).
Theorem 2.3.
Let $(\mathcal{S},\mathcal{X},\alpha)$ be an inverse semigroup action, where $\mathcal{X}$ is a compact Hausdorff space. The reduced crossed product $C_{0}(\mathcal{X})\rtimes_{r}\mathcal{S}$ is isomorphic to the reduced groupoid C-algebra $\textup{C}^{*}_{r}(\mathcal{S}\ltimes_{\alpha}\mathcal{X})$.
3. The Stone-Čech transformation groupoid of an inverse semigroup
Throughout this section, we identify $\beta\mathcal{S}$ with the space of ultrafilters on the discrete space $\mathcal{S}$. Given $A\subseteq\mathcal{S}$ we let
$U_{A}:=\{\xi\in\beta\mathcal{S}:A\in\xi\}.$
Sets of this form generate the topology on $\beta\mathcal{S}$. Each element $s\in\mathcal{S}$ determines a principal ultrafilter $\{A\subseteq\mathcal{S}:s\in A\}=\{s\}^{\subseteq}$; so the canonical embedding described in Example 2.2 is given by $\iota(s)=\{s\}^{\subseteq}$. In what follows we will drop the $\iota$ and consider $\mathcal{S}$ as a subset of $\beta\mathcal{S}$. We then have that $\overline{\mathcal{S}}=\beta\mathcal{S}$. Furthermore, it is a straightforward exercise to show that
$U_{A}=\overline{A},$
that is, the basic open set $U_{A}$ is the closure of $A$ in $\beta\mathcal{S}$.
Since $\mathcal{S}$ is a discrete space, there is a canonical homeomorphism $f:\overline{A}\to\beta A$. With ultrafilters, this map is defined by $f(\xi)=\{B\in\xi:B\subseteq A\}$.
Theorem 3.1.
The reduced groupoid C-algebra $\textup{C}^{*}_{r}(\mathcal{S}\ltimes\beta\mathcal{S})$ is isomorphic to the reduced crossed product $\ell^{\infty}(\mathcal{S})\rtimes_{r}\mathcal{S}$. In particular, $\mathcal{S}\ltimes\beta\mathcal{S}$ is amenable if and only if $\textup{C}^{*}_{r}(\mathcal{S})$ is exact.
Proof.
[…]
In what remains of this paper we study the action of $\mathcal{S}$ on its Stone-Čech compactification and properties of its transformation groupoid $\mathcal{S}\ltimes\beta\mathcal{S}$. In [EP16], Exel and Pardo proved many results about such actions and groupoids, so we recall some terminology and facts from there.
For $s\in\mathcal{S}$ we let
$\mathcal{J}_{s}$ $:=\{e\in E_{\mathcal{S}}:se=e\},$
$F_{s}$ $:=\{x\in\overline{s^{*}\mathcal{S}}:\lambda_{s}(x)=x\},$
$TF_{s}$ $:=\bigcup_{e\in\mathcal{J}_{s}}\overline{e\mathcal{S}}.$
Elements of $F_{s}$ are the fixed points of $s$ and the elements of $TF_{s}$ are the trivially fixed points of $s$; this is the terminology given in [EP16].
An element $[s,x]$ is in the unit space of $\mathcal{S}\ltimes\beta\mathcal{S}$ if and only if $[s,x]=[e,x]$ for some idempotent $e$, so we identify the unit space of $\mathcal{S}\ltimes\beta\mathcal{S}$ with $\beta\mathcal{S}$ via $e,x]\mapsto x$. If $[s,x]=[e,x]$ for $e\in E_{\mathcal{S}}$, there is an idempotent $f$ with $x\in\overline{f\mathcal{S}}$ with $sf=ef$. Thus $sfe=efe=fe$ and so $fe\in\mathcal{J}_{s}$. We also have $[s,x]=[e,x]=[ef,x]$.
It is clear that $x\in F_{s}$ if and only if $\lambda_{s}(x)=\mathbf{r}([s,x])=\mathbf{d}([s,x])=x$. In summary, we have observed the following lemma:
Lemma 3.2.
Let $\mathcal{S}$ be an inverse semigroup. If $s\in\mathcal{S}$, then
  1. (1)
    $x\in F_{s}\iff[s,x]\in\textup{Iso}(\mathcal{S}\ltimes\beta\mathcal{S})$;
  2. (2)
    $x\in TF_{s}\iff[s,x]\in(\mathcal{S}\ltimes\beta\mathcal{S})^{(0)}$;
  3. (3)
    $TF_{s}\subseteq F_{s}$.
A key dynamical fact for our purposes is the following consequence of Frolík:
Lemma 3.3.
Let $s\in\mathcal{S}$. If $\lambda_{s}\xi=\xi$ for some $\xi\in\overline{s^{*}\mathcal{S}}$, then there is a subset $A\subseteq s^{*}\mathcal{S}$ such that $\lambda_{s}\eta=\eta$, for all $\eta\in\overline{A}$. In particular, $\textup{Iso}(\mathcal{S}\ltimes\beta\mathcal{S})$ is open.
Proof.
The result is plain if $\xi$ is a principal ultrafilter, so we may assume that $\xi$ is a free ultrafilter (and therefore that $s^{*}\mathcal{S}$ is infinite.) Let $B\subset s^{*}\mathcal{S}$ be an infinite subset with infinite complement. Either $B\in\xi$ or $s^{*}\mathcal{S}\backslash B\in\xi$; without loss of generality, take $B\in\xi$. Because $\lambda_{s}:s^{*}\mathcal{S}\to s\mathcal{S}$ is a bijection, $sB\subset s\mathcal{S}$ is also an infinite subset with infinite complement. Because $\B\mathcal{S}\backslash$ B and $\sB\mathcal{S}\backslash$ sB are both countably infinite, we may extend the map $\lambda_{s}:B\to s^{*}B$ to a homeomorphism $f:\beta\mathcal{S}\to\beta\mathcal{S}$ with the property that $f(\xi)=\xi$. By Frolík’s theorem ([Frolik1971], Theorem 3.1), there is a basic open neighbourhood $U=\overline{C}$ of $\xi$ ($C\subseteq\mathcal{S}$) such that $f(\eta)=\eta$ for all $\eta\in U$. The result follows after taking $A=B\cap C$. ∎
We can now derive a nice form for the fixed point set of a given $s\in\mathcal{S}$.
Lemma 3.4.
If $\mathcal{S}$ is an inverse semigroup and $s\in\mathcal{S}$, then
(1) $F_{s}=\overline{\bigcup_{e\in\mathcal{J}_{s}}e\mathcal{S}}.$
Proof.
Suppose that there is some $\xi\in\overline{s^{*}\mathcal{S}}$ such that $\lambda_{s}\xi=\xi$. By Lemma 3.3, there exists $A\subseteq s^{*}\mathcal{S}$ such that $U_{A}=\overline{A}$ is fixed by $\lambda_{s}$. In particular, we have $sa=a$, for all $a\in A$, so $saa^{*}=aa^{*}$ for all $a\in A$. Said differently, $aa^{*}\in\mathcal{J}_{s}$, for all $a\in A$. Thus,
$A\subseteq\bigcup_{a\in A}a\mathcal{S}=\bigcup_{a\in A}aa^{*}\mathcal{S}\subseteq\bigcup_{e\in\mathcal{J}}e\mathcal{S}\implies\bigcup_{e\in\mathcal{J}}e\mathcal{S}\in\xi,$
because $\xi$ is a filter and filters are upwards closed. Thus $\xi\in\overline{\bigcup_{e\in\mathcal{J}}e\mathcal{S}}$, and so $F_{s}\subseteq\overline{\bigcup_{e\in\mathcal{J}_{s}}e\mathcal{S}}$. Reverse containment follows from the continuity of $\lambda_{s}$, completing the proof. ∎
Definition 3.5.
Let $\mathcal{G}$ be a groupoid. We say that
  1. (1)
    $\mathcal{G}$ is principal if $g\in\mathcal{G}$ and $s(g)=r(g)$ implies $g\in\mathcal{G}^{(0)}$,
  2. (2)
[…]
  3. (3)
    $\mathcal{G}$ is essentially principal if $\{x\in\mathcal{G}^{(0)}:\mathcal{G}_{x}^{x}=\{x\}\}$ is dense in $\mathcal{G}^{(0)}$.
Proposition 3.6.
Let $s\in\mathcal{S}$ and suppose $s,x]\in\textup{Iso}(\mathcal{S}\ltimes\beta\mathcal{S})$ for some principal filter $x\in s^{*}\mathcal{S}$. Then $s,x]\in(\mathcal{S}\ltimes\beta\mathcal{S})^{(0)}$. In particular, $\mathcal{S}\ltimes\beta\mathcal{S}$ is always essentially principal.
Proof.
We have that $sx=x$ and so $sxx^{*}=xx^{*}$. Since $x\in xS=xx^{*}\mathcal{S}$, we have that $xx^{*}\mathcal{S}\in x$. Thus $s,x]=[sxx^{*},x]=[xx^{*},x]\in(\mathcal{S}\ltimes\beta\mathcal{S})^{(0)}$. Because principal filters are dense in $(\mathcal{S}\ltimes\beta\mathcal{S})^{(0)}$, $\mathcal{S}\ltimes\beta\mathcal{S}$ is essentially principal. ∎
Theorem 3.7.
For an inverse semigroup $\mathcal{S}$, the following conditions are equivalent:
  1. (1)
    $\mathcal{S}\ltimes\beta\mathcal{S}$ is Hausdorff;
  2. (2)
    $\mathcal{S}\ltimes\beta\mathcal{S}$ is principal;
  3. (3)
    $\mathcal{S}\ltimes\beta\mathcal{S}$ is effective.
Proof.
$(1)\implies(2)$ By [EP16], $\mathcal{S}\ltimes\beta\mathcal{S}$ is Hausdorff if and only if $TF_{s}=\bigcup_{e\in\mathcal{J}_{s}}\overline{e\mathcal{S}}$ is closed relative to $\overline{s^{*}\mathcal{S}}$. Lemma 3.4 says that $F_{s}$ is the smallest closed set that contains $e\mathcal{S}$ for all $e\in\mathcal{J}_{s}$, so $F_{s}\subseteq TF_{s}$, and hence $F_{s}=TF_{s}$. Because every fixed point is trivially fixed, it follows from Lemma 3.2 that $\mathcal{S}\ltimes\beta\mathcal{S}$ is principal.
$(2)\implies(1)$ if $\mathcal{G}$ is not Hausdorff, we can find a net $(\gamma_{i})$ in $\mathcal{G}$ with two distinct limits $\gamma,\gamma^{\prime}$. Because the range and source maps are continuous, we have $r(\gamma)=r(\gamma^{\prime})$ and $s(\gamma)=s(\gamma^{\prime})$. Because $\gamma\neq\gamma^{\prime}$, $\mathcal{G}$ is not principal, and so by the contrapositive we are done.
$(2)\iff(3)$ follows from Lemma 3.3. ∎
In light of Theorem 3.7, the key structural question one asks is when is $\mathcal{S}\ltimes\beta\mathcal{S}$ Hausdorff? Below is a criterion that produces examples.
Theorem 3.8.
Let $\mathcal{S}$ be an inverse semigroup. Then the following conditions are equivalent:
  1. (1)
    $\mathcal{S}\ltimes\beta\mathcal{S}$ is Hausdorff;
  2. (2)
    for every $s\in\mathcal{S}$, there exists a finite set $F\subseteq\mathcal{J}_{s}$ such that $\bigcup_{f\in F}f\mathcal{S}=\bigcup_{e\in\mathcal{J}_{s}}e\mathcal{S}$;
  3. (3)
    for every $s\in\mathcal{S}$, there exists a finite set $F\subseteq\mathcal{J}_{s}$ such that $\mathcal{J}_{s}=F^{\geqslant}$.
Proof.
$(1)\implies(2)$ By Theorem 3.7, if $\mathcal{S}\ltimes\beta\mathcal{S}$ is Hausdorff, then $TF_{s}=F_{s}=\bigcup_{e\in\mathcal{J}_{s}}\overline{e\mathcal{S}}$ for all $s\in\mathcal{S}$. This set is compact and each $\overline{e\mathcal{S}}$ is open, so there is a finite $F\subseteq\mathcal{J}_{s}$ such that
$\overline{\bigcup_{e\in\mathcal{J}_{s}}e\mathcal{S}}=\bigcup_{e\in\mathcal{J}_{s}}\overline{e\mathcal{S}}=\bigcup_{f\in F}\overline{f\mathcal{S}}=\overline{\bigcup_{f\in F}f\mathcal{S}},$
where the last equality follows as $F$ is finite. Because elements of $\mathcal{S}$ are isolated in $\beta\mathcal{S}$, this implies $\bigcup_{f\in F}f\mathcal{S}=\bigcup_{e\in\mathcal{J}_{s}}e\mathcal{S}$.
$(2)\implies(3)$ For each $e\in\mathcal{J}_{s}$, there exists $f\in F$ such that $e\in f\mathcal{S}$ so we can write $e=ft$ for some $t\in\mathcal{S}$. Thus $fe=fft=ft=e$, implying $e\leqslant f$. Conversely, if $e\leqslant f\in F\subseteq\mathcal{J}_{s}$, $se=sfe=fe=e$ and so $e\in\mathcal{J}_{s}$.
$(3)\implies(2)$ If $e\in\mathcal{J}_{s}$, then by assumption, $e\leqslant f$ for some $f\in F$; so $e\in f\mathcal{S}$. Thus, we have $\bigcup_{f\in F}f\mathcal{S}=\bigcup_{e\in\mathcal{J}_{s}}e\mathcal{S}$.
$(2)\implies(1)$ We have
$F_{s}=\overline{\bigcup_{e\in\mathcal{J}_{s}}e\mathcal{S}}=\overline{\bigcup_{f\in F}f\mathcal{S}}=\bigcup_{f\in F}\overline{f\mathcal{S}}\subseteq\bigcup_{e\in\mathcal{J}_{s}}\overline{e\mathcal{S}}=TF_{s},$
implying $F_{s}=TF_{s}$. By Theorem 3.7, $\mathcal{S}\ltimes\beta\mathcal{S}$ is Hausdorff. ∎
Two elements $s,t\in\mathcal{S}$ are called compatible if both $s^{*}t$ and $st^{*}$ are idempotents. Notice that idempotents are always compatible. The inverse semigroup $\mathcal{S}$ is called complete if whenever $A\subseteq\mathcal{S}$ is a set of pairwise compatible elements, then $A$ has a least upper bound in $\mathcal{S}$ (in the natural partial order); in this case the least upper bound is denoted $\bigvee_{a\in A}a$. An inverse semigroup is called infinitely distributive if whenever $\bigvee_{a\in A}a\in\mathcal{S}$ and $s\in\mathcal{S}$, we have
$s\left(\bigvee_{a\in A}a\right)=\bigvee_{a\in A}sa\hskip 28.45274pt\text{and}\hskip 28.45274pt\left(\bigvee_{a\in A}a\right)s=\bigvee_{a\in A}as.$
An inverse semigroup is an abstract pseudogroup if it is a complete, infinitely distributive inverse monoid (see [lawson]). The symmetric inverse monoid $\mathcal{I}(X)$ is an abstract pseudogroup. An inverse semigroup with zero $\mathcal{S}$ is $E^{*}$-unitary if $s\in\mathcal{S}$ and $\mathcal{J}_{s}\neq\{0\}$ implies $s\in E_{S}$. Graph inverse semigroups and free inverse semigroups are $E^{*}$-unitary [L]. The next result is an immediate consequence of Theorem 3.8.
Corollary 3.9.
If $\mathcal{S}$ is an abstract pseudogroup or an $E^{*}$-unitary inverse semigroup, then $\mathcal{S}\ltimes\beta\mathcal{S}$ is Hausdorff.
Proof.
If $\mathcal{S}$ is an abstract pseudogroup and $s\in\mathcal{S}$, then $\mathcal{J}_{s}$ is pairwise compatible and hence $e:=\bigvee_{a\in\mathcal{J}_{s}}a\in\mathcal{S}$. We also have that $se=\bigvee_{a\in\mathcal{J}_{s}}sa=\bigvee_{a\in\mathcal{J}_{s}}a=e$ by distributivity, so $e\in\mathcal{J}_{s}$. The finite set $F=\{e\}$ satisfies Theorem 3.8(3).
Every transformation groupoid of an $E^{*}$-unitary inverse semigroup is Hausdorff by [EP16]. Alternatively, observe that if $\mathcal{J}_{s}$ is nonzero, then $s\in\mathcal{J}_{s}$, and thus $\mathcal{J}_{s}=\{s\}^{\geqslant}$. ∎
The main result of [starling] says that the action of $\mathcal{S}$ on its Exel spectrum is amenable if and only if every action of $\mathcal{S}$ is amenable. Our last result has a similar flavour, and provides some clear obstructions to the Hausdorffness of $\mathcal{S}\ltimes\beta\mathcal{S}$.
Definition 3.10.
Say that an action $(\mathcal{S},\mathcal{X},\alpha)$ is clopen if $D^{\alpha}_{e}$ is clopen for each $e\in E_{\mathcal{S}}$.
Corollary 3.11.
For an inverse semigroup $\mathcal{S}$, the following conditions are equivalent:
  1. (1)
    $\mathcal{S}\ltimes\beta\mathcal{S}$ is Hausdorff;
  2. (2)
    $\mathcal{S}\ltimes\mathcal{X}$ is Hausdorff for every clopen action $(\mathcal{S},\mathcal{X},\alpha)$.
Proof.
[…]
Let $s\in\mathcal{S}$. By assumption, $\alpha_{s}:D^{\alpha}_{s^{*}s}\to D^{\alpha}_{ss^{*}}$ a homeomorphism between clopen sets. Writing $TF^{\alpha}_{s}$ for the trivially fixed points for the action $\alpha$, we have
$TF^{\alpha}_{s}=\bigcup_{e\in\mathcal{J}_{s}}D^{\alpha}_{e}.$
Applying Theorem 3.8, we can find a finite set $F\subseteq\mathcal{J}_{s}$ such that, for all $e\in\mathcal{J}_{s}$, there exists $f\in F$ with $e\leqslant f$. Since $D_{e}^{\alpha}\subseteq D_{f}^{\alpha}$ whenever $e\leqslant f$, we have
$TF^{\alpha}_{s}=\bigcup_{e\in\mathcal{J}_{s}}D^{\alpha}_{e}=\bigcup_{f\in F}D^{\alpha}_{f}.$
By assumption, each $D^{\alpha}_{f}$ is closed. Since $F$ is finite, $TF^{\alpha}_{s}$ is closed, so by [EP16, Theorem 3.15], $\mathcal{S}\ltimes\mathcal{X}$ is Hausdorff. ∎
Corollary 3.12.
If $\mathcal{S}\ltimes\beta\mathcal{S}$ is Hausdorff, then Exel’s tight groupoid of $\mathcal{S}$, Exel’s universal groupoid of $\mathcal{S}$, and Paterson’s universal groupoid of $\mathcal{S}$ are all Hausdorff.
Negating Corollary 3.11 then shows that we can find many examples where $\mathcal{S}\ltimes\beta\mathcal{S}$ is not Hausdorff (and hence not principal), namely any example where the tight groupoid is not Hausdorff. See [CEPSS, Example 5.6] for an example arising from the Grigorchuk group.
Question: does Corollary 3.11 hold for non-clopen actions?